# Toward Sufficient Statistical Power in Algorithmic Bias Assessment: A Test for ABROCA

### v1

@@ Abstract
Algorithmic bias is a pressing concern in educational data mining (EDM), as it risks amplifying inequities in learning outcomes. The Area Between ROC Curves (ABROCA) metric is frequently used to measure discrepancies in model performance across demographic groups to quantify overall model fairness. However, its skewed distribution—especially when class or group imbalances exist—makes significance testing challenging. This study investigates ABROCA’s distributional properties and contributes robust methods for its significance testing. Specifically, we address (1) whether ABROCA follows any known distribution, (2) how to reliably test for algorithmic bias using ABROCA, and (3) the statistical power achievable with ABROCA-based bias assessments under typical EDM sample specifications. Simulation results confirm that ABROCA does not match standard distributions, including those suited to accommodate skewness. We propose nonparametric randomization tests for ABROCA and demonstrate that reliably detecting bias with ABROCA requires large sample sizes or substantial effect sizes, particularly in imbalanced settings. Findings suggest that ABROCA-based bias evaluation based on sample sizes common in EDM tends to be underpowered, undermining the reliability of conclusions about model fairness. By offering open-source code to simulate power and statistically test ABROCA, this paper aims to foster more reliable statistical testing in EDM research. It supports broader efforts toward replicability and equity in educational modeling.
@@ 1 Introduction
Algorithmic bias is a critical concern in educational data mining (EDM), as it can undermine the effectiveness of key EDM applications. Past research has studied algorithmic bias in knowledge tracing [[22]], grade prediction [[10], [5], [17]], and at-risk prediction [[4], [11]]. Despite differences in definitions [[1]], algorithmic bias generally refers to systematic unfairness or unequal treatment caused by algorithms, often disadvantaging or favoring groups based on immutable traits such as race, gender, or socioeconomic status [[7]]. In this context, efforts to measure algorithmic fairness in EDM models have focused on metrics that quantify performance differences between student groups (e.g., by race or gender). A model is typically considered fair if its predictive performance is equivalent or comparable across groups.
@@
However, like all statistical metrics, measures of algorithmic bias are subject to random fluctuation through sampling. A critical yet underexplored dimension in algorithmic bias research is the role of statistical power in bias detection. Statistical power represents the probability of correctly rejecting the null hypothesis—concluding a model is biased when it truly is. Inferential conclusions about model fairness are vulnerable to both Type I (false positive) and Type II (false negative) errors. For example, insufficient statistical power might lead to the erroneous conclusion that a model is fair, masking genuine disparities in performance among demographic groups (false negative). Conversely, a model may be incorrectly deemed biased when it is not (false positive). While considerations of statistical power are widely recognized in other EDM applications [[9]], prior research reporting that models are fair have typically not considered statistical power [[10], [22], [3]]. Furthermore, systematic investigations into how sample sizes, effect sizes, and design factors influence the reliability of bias detection remain scarce. This gap highlights an urgent need to incorporate robust power analysis and reporting standards to ensure fairness claims—or their absence—are replicable, meaning they are generalizable to other samples [[9]].
@@
Among the various metrics for assessing algorithmic bias, ABROCA is an illustrative example that underscores the importance of statistical power considerations. ABROCA measures discrepancies in classification performance between demographic groups by comparing the area under their respective receiver operating characteristic (ROC) curves. Larger ABROCA values indicate greater performance disparities, signaling substantial algorithmic bias. However, recent studies have highlighted the skewed distribution of ABROCA, particularly in datasets with class or group imbalances [[3]]. In other words, large ABROCA values may arise purely by chance, even when no population-level bias or differences in ROC curves exist. This can lead to false-positive conclusions about a model’s bias based on observed ABROCA values, driven by random sampling variability.
@@
By addressing these questions, we provide practical guidance for researchers on selecting statistical methods to test bias using ABROCA while minimizing false positives and ensuring reliable claims of fairness. To promote reproducibility, we contribute open-source code for simulating statistical power and testing ABROCA under varying conditions. [^1] https://anonymous.4open.science/r/bias-power This work enhances the rigor of algorithmic fairness research in EDM, contributing to emerging fairness evaluation methods.
@@ 2.3 The ABROCA Metric
ABROCA has become a widely used metric for assessing model fairness. Gardner et al. [[8]] introduced ABROCA as a method to capture performance differences between classifiers by examining precision-recall tradeoff thresholds. Unlike metrics that focus solely on AUC ROC values, ABROCA evaluates regions of the ROC curve where a model may exhibit unequal performance across demographic groups, providing a more nuanced view of model fairness.
@@
These studies collectively demonstrate ABROCA’s versatility in measuring fairness across a range of contexts in EDM. They also highlight the importance of understanding the metric’s statistical properties, as ABROCA is often used to assess a classifier’s overall algorithmic fairness, including through methods of statistical comparison and significance testing.
@@ 3.1 RQ1: ABROCA’s Distribution
To investigate whether ABROCA follows a known distribution, we generated a simulated dataset of ABROCA values across various conditions, including variations in sample size, class imbalance ratios, and outcome imbalance ratios, following Borchers and Baker [[3]]. This investigation was conducted under the null hypothesis that both groups share the same AUC values and that the model is equally accurate for both groups. Simulations were performed under two experimental conditions: (1) balanced data, where both groups and outcome classes were of equal size, and (2) imbalanced data, with a 90% imbalance ratio. Each experiment used 5,000 observations.
@@ 3.2 RQ2: Simulating Statistical Power
Statistical power increases with sample size and more precise estimates [[20]]. To evaluate power in the specific context of algorithmic bias analysis with ABROCA, it is necessary to develop statistical tests and simulation procedures. This study introduces a statistical test for ABROCA, enabling power estimation for common sample sizes in the field (see Section 3.3 and Algorithm 1).
@@
The results provide insights into the ability of ABROCA-based statistical tests to detect true differences in bias. We evaluate test set sizes between 100 and 2,000 samples total, which is typical in educational data mining research zambrano2024investigating ; vsvabensky2024evaluating .
@@ 3.3 RQ3: Statistical Power and Imbalance
We propose a nonparametric randomization test to assess the significance of a given ABROCA value (Algorithm 1). Randomization tests are a common method for nonparametric significance testing zhang2023randomization . For the present study’s context, the test premise is that, for a classifier free from algorithmic bias, the group membership of observations should not influence the magnitude of the ABROCA metric. This test involves randomly shuffling group memberships between the two groups, defining the ABROCA metric, and generating a null distribution of ABROCA values. A p-value is then calculated by comparing the observed ABROCA value (based on the actual group assignment) to the null distribution derived from random group assignments. This allows for a comparison of the observed ABROCA against the null hypothesis, producing a p-value that represents the probability of observing an ABROCA value as extreme as the actual one under the assumption of no group disparity (Figure 1).
@@ 4.1 RQ1: ABROCA’s Distribution
Figure 2: Comparison of Q-Q plots for imbalanced and balanced null hypotheses based on a sample size of 5,000. The plots illustrate the deviations of ABROCA sample quantiles from theoretical quantiles for various tested distributions, including Weibull, normal, $t$-distribution, and $F$-distribution. These comparisons help evaluate which distributions best approximate ABROCA under different class balance conditions.
In summary, ABROCA does not conform to any standard known distribution. Among the flexible and skew-accommodating distributions tested, the Weibull distribution fits the data best (but still significantly deviates from the theoretical distribution). As shown in the Q-Q plots (Figure 2), the Weibull distribution exhibits deviations at the extreme distribution ends in both balanced and imbalanced scenarios. This breakdown at the tails suggests that even under ideal conditions, the Weibull distribution may not be suitable for accurately modeling ABROCA.
@@ 5 Discussion
Our simulations confirmed that ABROCA does not conform to any standard distribution (including normal, $t$-, $F$-, and Weibull). Even though Weibull provided a closer fit than the other candidates, it still exhibited systematic deviations, particularly in the distribution tails. Consequently, relying on off-the-shelf parametric tests (e.g., $t$-tests) for ABROCA can be problematic, as standard distributional assumptions are not met. The distributional skewness and heavy tails reinforce that ABROCA requires more specialized inference methods. We propose randomization tests as a nonparametric alternative to testing whether an obtained ABROCA value is significantly above chance, an issue that recent research has identified as a key issue in the widespread use of ABROCA, given its heavy skew borchers2025abroca . Researchers and practitioners may use the open-source code presented here to reproduce the randomized testing procedure, including sample size planning through statistical power simulation converging within a few hours on contemporary standard computers.
@@ 5.2 Limitations and Future Work
There are several limitations to the present study. First, we focused on holdout test sets rather than the common cross-validation procedure in EDM. Leave-one-out cross-validation is generally asymptotically equivalent to comparing model $AIC$ shao1993linear , an alternative criterion to determining whether two models or theories (e.g., that of a biased or unbiased model) describe empirical data better. More research is needed to describe the reliability of algorithmic bias assessments under such research frameworks. Second, like many simulation studies, our results rely on synthetic data-generating processes, which may not capture all the complexities of real-world educational datasets (e.g., missing data, high-dimensional feature spaces, or intricate dependencies) Kaser2024simulated . Future research could examine how these intricacies affect bias detection and ABROCA’s distribution. Third, we did not propose a specific minimal (practically) relevant effect size for ABROCA. Determining this threshold is domain- and context-dependent, requiring input from practitioners and policy-makers regarding what magnitude of bias is ethically or educationally meaningful. Deliberation of minimally relevant effect sizes during the study design phase is foundational to sample size planning using statistical power simulations wickelmaier2021simulating .
@@ 6 Conclusion
This study examined the distribution of the ABROCA metric and introduced a nonparametric randomization test to assess its significance for algorithmic bias detection. Our simulations confirmed that ABROCA’s skewness and heavy-tailed behavior make the metric ill-suited to standard parametric tests. We demonstrated that in many realistic scenarios—particularly those involving modest sample sizes and imbalanced group or outcome class distributions—ABROCA-based bias analyses can suffer from low statistical power, potentially leading to inconclusive results about fairness. Consequently, studies might fail to detect genuine model bias or might incorrectly infer bias where none exists due to random chance.
By providing open-source code to reproduce our randomization-based significance tests and power analyses, we enable other researchers to evaluate ABROCA in diverse contexts. Our findings underscore the need for a cautious interpretation of ABROCA values: the absence of evidence for bias is not necessarily evidence of fairness if the analysis is underpowered. In turn, researchers should plan for sufficient sample sizes or adopt data pooling and oversampling strategies when working with imbalanced datasets. We hope this work encourages the broader adoption of robust statistical frameworks in algorithmic bias assessment, ultimately supporting more reliable model evaluations and equitable model application outcomes.

### v2

@@ Abstract
Algorithmic bias is a pressing concern in educational data mining (EDM), as it risks amplifying inequities in learning outcomes. The Area Between ROC Curves (ABROCA) metric is frequently used to measure discrepancies in model performance across demographic groups to quantify overall model fairness. However, its skewed distribution–especially when class or group imbalances exist–makes significance testing challenging. This study investigates ABROCA’s distributional properties and contributes robust methods for its significance testing. Specifically, we address (1) whether ABROCA follows any known distribution, (2) how to reliably test for algorithmic bias using ABROCA, and (3) the statistical power achievable with ABROCA-based bias assessments under typical EDM sample specifications. Simulation results confirm that ABROCA does not match standard distributions, including those suited to accommodate skewness. We propose nonparametric randomization tests for ABROCA and demonstrate that reliably detecting bias with ABROCA requires large sample sizes or substantial effect sizes, particularly in imbalanced settings. Findings suggest that ABROCA-based bias evaluations based on sample sizes common in EDM tend to be underpowered, undermining the reliability of conclusions about model fairness. By offering open-source code to simulate power and statistically test ABROCA, this paper aims to foster more reliable statistical testing in EDM research. It supports broader efforts toward replicability and equity in educational modeling.
@@ 1 Introduction
Algorithmic bias is a critical concern in educational data mining (EDM), as it can undermine the effectiveness of key EDM applications. Past research has studied algorithmic bias in knowledge tracing [[26]], grade prediction [[14], [8], [21]], and at-risk prediction [[7], [15]]. Despite differences in definitions [[2]], algorithmic bias generally refers to systematic unfairness or unequal treatment caused by algorithms, often disadvantaging or favoring certain individuals or groups of individuals in favor of others [[10]]. In this context, efforts to measure algorithmic fairness in EDM models have often focused on metrics that quantify performance differences between student groups based on demographic attributes (e.g., by race or gender). A model is typically considered fair if its predictive performance is equivalent or comparable across groups.
@@
However, like all statistical metrics, measures of algorithmic bias are subject to random fluctuation through sampling. A critical yet underexplored dimension in algorithmic bias research is the role of statistical power in bias detection. Statistical power represents the probability of correctly rejecting the null hypothesis—concluding a model is biased when it truly is. Inferential conclusions about model fairness are vulnerable to both Type I (false positive) and Type II (false negative) errors. For example, insufficient statistical power might lead to the erroneous conclusion that a model is fair, masking genuine disparities in performance among demographic groups (false negative). Conversely, a model may be incorrectly deemed biased when it is not (false positive). While considerations of statistical power are widely recognized in other EDM applications [[12]], prior research reporting that models are fair have typically not considered statistical power [[14], [26], [4]]. Furthermore, systematic investigations into how sample sizes, effect sizes, and design factors influence the reliability of bias detection remain scarce. One exception is Choi et al. [[6]], who applied bootstrapping to a dropout prediction data set to quantify uncertainty in bias metrics through confidence intervals. This gap highlights an urgent need to incorporate robust power analysis and reporting standards to ensure fairness claims—or their absence—are replicable, meaning they are generalizable to other samples [[12]].
@@
Among the various metrics for assessing algorithmic bias, ABROCA is an illustrative example that underscores the importance of statistical power considerations. ABROCA measures discrepancies in classification performance between demographic groups by comparing the area under their respective receiver operating characteristic (ROC) curves. Larger ABROCA values indicate greater performance disparities, signaling substantial algorithmic bias. However, recent studies have highlighted the skewed distribution of ABROCA, particularly in datasets with class or group imbalances [[4]]. In other words, large ABROCA values may arise purely by chance, even when no population-level bias or differences in ROC curves exist. This can lead to false-positive conclusions about a model’s bias based on observed ABROCA values driven by random sampling variability.
@@
By addressing these questions, we provide practical guidance for researchers on selecting statistical methods to test bias using ABROCA while minimizing false positives and ensuring reliable claims of fairness. To promote reproducibility, we contribute open-source code for simulating statistical power and testing ABROCA under varying conditions. [^1] https://github.com/conradborchers/abroca-test-power This work enhances the rigor of algorithmic fairness research in EDM, contributing to emerging fairness evaluation methods.
@@ 2.3 The ABROCA Metric
A widely used measure for evaluating binary classifiers is the area under the ROC curve (AUC) metric [[5]]. The AUC metric offers an intuitive interpretation (the probability that the classifier correctly distinguishes a positive instance from a negative one) and reliable performance assessment for imbalanced class distributions [[13]]. Its robustness stems from evaluating the true positive rate (TPR) and false positive rate (FPR) across all classification thresholds from 0 to 1. The ROC curve plots TPR against FPR at varying thresholds, and the AUC quantifies the area beneath this curve. A higher AUC indicates better performance, with 0.5 representing chance-level prediction and 1 denoting perfect classification.
ABROCA is a widely adopted fairness metric that measures disparities in classifier performance across demographic subgroups. Introduced by Gardner et al. [[11]], it quantifies biases along the decision-threshold spectrum by computing the area between subgroup-specific ROC curves. Unlike traditional AUC-based comparisons, which assess overall performance, ABROCA identifies localized disparities—revealing where a model may favor one group over another at specific thresholds. This granular approach makes ABROCA particularly valuable for detecting nuanced fairness issues that aggregate metrics might overlook.
@@
These studies collectively demonstrate ABROCA’s versatility in measuring fairness across various contexts in EDM. They also highlight the importance of understanding the metric’s statistical properties, as ABROCA is often used to assess a classifier’s overall algorithmic fairness, including through methods of statistical comparison and significance testing.
@@ 3.1 RQ1: ABROCA’s Distribution
To investigate whether ABROCA follows a known distribution, we generated a simulated dataset of ABROCA values across various conditions, including variations in sample size, class imbalance ratios, and outcome imbalance ratios, following Borchers and Baker [[4]]. This investigation was conducted under the null hypothesis that both groups share the same AUC values and that the model is equally accurate for both groups. Simulations were performed under two experimental conditions: (1) balanced data, where both groups and outcome classes were of equal size, and (2) imbalanced data, with a 90% imbalance ratio for both groups and outcomes. For this distributional analysis, 5,000 observations were simulated from each distribution.
@@ 3.2 RQ2: Simulating Statistical Power
Statistical power increases with sample size and more precise estimates [[24]]. Statistical tests and simulation procedures must be developed to evaluate power in the specific context of algorithmic bias analysis with ABROCA. This study introduces a statistical test for ABROCA, enabling power estimation for common sample sizes in the field (see Section 3.3 and Algorithm 1).
@@
The results provide insights into the ability of ABROCA-based statistical tests to detect true differences in bias. We evaluate test set sizes between 100 and 2,000 samples, which is typical in educational data mining research zambrano2024investigating ; vsvabensky2024evaluating .
@@ 3.3 RQ3: Statistical Power and Imbalance
[…]
A density plot titled ’Density’ showing the distribution of the ABROCA statistic. The x-axis is labeled ’ABROCA Statistic Value’ and includes values such as 0.025, 0.050, 0.075, and 0.100. The plot features two labeled components: ’Threshold’ and ’Observed Statistic,’ indicating key points within the distribution.
@@ 4.1 RQ1: ABROCA’s Distribution
[…]
A series of four Q-Q (Quantile-Quantile) plots comparing sample distributions to theoretical distributions for imbalanced data. Each plot is labeled with its type: Weibull, t-distribution, Normal, and F-distribution. Below each plot, the Kolmogorov-Smirnov (KS) statistic and p-value are displayed, indicating the goodness-of-fit (e.g., ’KS Statistic: 0.076, p < .001’ for Weibull). The x-axes represent theoretical quantiles, while the y-axes show sample quantiles. The KS statistics suggest significant deviations from the tested distributions, with the largest deviation occurring for the F-distribution (KS = 0.554). The plots are used to assess how well the sample data matches each theoretical model.
@@
[…]
A series of four Q-Q (Quantile-Quantile) plots comparing sample distributions to theoretical distributions for balanced data. Each plot is labeled with its type: Weibull, t-distribution, Normal, and F-distribution. Below each plot, the Kolmogorov-Smirnov (KS) statistic and p-value are displayed, indicating the goodness-of-fit (e.g., ’KS Statistic: 0.091, p < .001’ for Weibull). The x-axes represent theoretical quantiles, while the y-axes show sample quantiles. The KS statistics suggest significant deviations from the tested distributions, with the largest deviation occurring for the F-distribution (KS = 0.689). The plots are used to assess how well the sample data matches each theoretical model.
@@
Figure 2: Comparison of Q-Q plots for imbalanced and balanced null hypotheses based on a sample size of 5,000. The plots illustrate the deviations of ABROCA sample quantiles from theoretical quantiles for various tested distributions, including Weibull, normal, $t$-distribution, and $F$-distribution. These comparisons help evaluate which distributions best approximate ABROCA under different class balance conditions. The red line represents the expected relationship for a distribution match.
In summary, ABROCA does not conform to any standard known distribution. Among the skew-accommodating and flexible distributions tested, the Weibull distribution fits the data best (but still significantly deviates from the theoretical distribution). As shown in the Q-Q plots (Figure 2), the Weibull distribution exhibits deviations at the extreme distribution ends in both balanced and imbalanced scenarios. This breakdown at the tails suggests that even under ideal conditions, the Weibull distribution may not be suitable for accurately modeling ABROCA.
@@ 4.2 RQ2: A Statistical Test for ABROCA
[…]
A line graph titled ’Effect Size (AUC Difference)’ showing the relationship between evaluation sample size (x-axis) and statistical power (y-axis, ranging from 0.0 to 1.0). The power curve illustrates how the probability of detecting an effect (AUC difference) increases with larger sample sizes. The graph is used to assess the required sample size for achieving desired statistical power (e.g., 0.8 or higher) in a study measuring AUC differences.
@@ 4.3 RQ3: Impact of Imbalance on Power
[…]
A power analysis graph titled ’Imbalance (AUC1 = 0.8, AUC2 = 0.7)’ showing the relationship between evaluation sample size (x-axis, ranging from 100 to 2000) and statistical power (y-axis). The plot compares four data balance or imbalance conditions labeled ’None’, ’Both’, ’Group’, and ’Label’, representing different sampling methods. The graph illustrates how power increases with larger sample sizes when detecting an AUC difference between two models (0.8 vs. 0.7). This visualization helps determine the required sample size to achieve adequate statistical power for fairness or performance comparisons.
@@ 5 Discussion
Simulations confirmed that ABROCA does not conform to any standard distribution (including normal, $t$-, $F$-, and Weibull). Even though Weibull provided a closer fit than the other candidates, it still exhibited systematic deviations, particularly in the distribution tails. Consequently, relying on off-the-shelf parametric tests (e.g., $t$-tests) for ABROCA can be problematic, as standard distributional assumptions are not met. The distributional skewness and heavy tails reinforce that ABROCA requires more specialized inference methods. We propose randomization tests as a nonparametric alternative to testing whether an obtained ABROCA value is significantly above chance, an issue that recent research has identified as a key issue in the widespread use of ABROCA, given its heavy skew borchers2025abroca . Researchers and practitioners may use the open-source code presented here to reproduce the randomized testing procedure, including sample size planning through statistical power simulation converging within a few hours on contemporary standard computers.
@@ 5.2 Limitations and Future Work
There are several limitations to the present study. First, we focused on holdout test sets rather than the common cross-validation procedure in EDM. Leave-one-out cross-validation is generally asymptotically equivalent to comparing model $AIC$ shao1993linear , an alternative criterion to determining whether two models or theories (e.g., that of a biased or unbiased model) describe empirical data better. More research is needed to describe the reliability of algorithmic bias assessments under such research frameworks. Second, like many simulation studies, our results rely on synthetic data-generating processes, which may not capture all the complexities of real-world educational datasets (e.g., missing data, high-dimensional feature spaces, or intricate dependencies) Kaser2024simulated . Future research could examine how these intricacies affect bias detection and ABROCA’s distribution using our open-source code. These applications may include setting simulation parameters according to real-world data sets or applying our testing procedure to data directly. Third, we did not propose a specific minimal (practically) relevant effect size for ABROCA. Determining this threshold is domain- and context-dependent, requiring input from practitioners and policy-makers regarding what magnitude of bias is ethically or educationally meaningful. Deliberation of minimally relevant effect sizes during the study design phase is foundational to sample size planning using statistical power simulations wickelmaier2021simulating . In our field, these deliberations could consider whether an intervention is fail-soft or high-stakes as well as its cost (where cost could be conceived as economic, ethical, or otherwise), especially in situations where it may be challenging to obtain high-precision prediction models baker2025difficulty .
@@ 6 Conclusion
This study examined the distribution of the ABROCA metric and introduced a nonparametric randomization test to assess its significance for algorithmic bias detection. Our simulations confirmed that ABROCA’s skewness and heavy-tailed behavior make the metric ill-suited to standard parametric tests. We demonstrated that in many realistic scenarios—particularly those involving modest sample sizes and imbalanced group or outcome class distributions—ABROCA-based bias analyses can suffer from low statistical power, potentially leading to inconclusive results about fairness. Consequently, studies might fail to detect genuine model bias or incorrectly infer bias where none exists due to random chance.
By providing open-source code to reproduce our randomization-based significance tests and power analyses, we enable other researchers to evaluate ABROCA in diverse contexts. Our findings underscore the need for a cautious interpretation of ABROCA values: the absence of evidence for bias is not necessarily evidence of fairness if the analysis is underpowered. Researchers should plan for sufficient sample sizes or adopt data pooling and oversampling strategies when working with imbalanced datasets. We hope this work encourages the broader adoption of robust statistical frameworks in algorithmic bias assessment, ultimately supporting more reliable model evaluations and equitable model application outcomes.
Acknowledgments
I thank Ryan S. Baker for pointing out the intricate relationship between cross-validation and statistical testing (see Section 5.2) and for earlier discussions that led to the inception of this research.